### v1

# Fine-Tuning Large Language Models to Translate: Will a Touch of Noisy Data in Misaligned Languages Suffice?

Abstract
Traditionally, success in multilingual machine translation can be attributed to three key factors in training data: large volume, diverse translation directions, and high quality. In the current practice of fine-tuning large language models (LLMs) for translation, we revisit the importance of all these factors. We find that LLMs display strong translation capability after being fine-tuned on as few as 32 training instances, and that fine-tuning on a single translation direction effectively enables LLMs to translate in multiple directions. However, the choice of direction is critical: fine-tuning LLMs with English on the target side can lead to task misinterpretation, which hinders translations into non-English languages. A similar problem arises when noise is introduced into the target side of parallel data, especially when the target language is well-represented in the LLM’s pre-training. In contrast, noise in an under-represented language has a less pronounced effect. Our findings suggest that attaining successful alignment hinges on teaching the model to maintain a “superficial” focus, thereby avoiding the learning of erroneous biases beyond translation.
Fine-Tuning Large Language Models to Translate:
Will a Touch of Noisy Data in Misaligned Languages Suffice?
Dawei Zhu1 Pinzhen Chen2 Miaoran Zhang1 Barry Haddow2 Xiaoyu Shen3 Dietrich Klakow1 1Saarland University, Saarland Informatics Campus 2University of Edinburgh 3Eastern Institute of Technology {dzhu, mzhang}@lsv.uni-saarland.de pinzhen.chen@ed.ac.uk
1 Introduction
Large language models (LLMs) have reached heights in various NLP tasks (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023). Supervised fine-tuning (SFT (Ouyang et al., 2022), also called instruction tuning or simply fine-tuning in some literature) further prepares these models for better generalization and reliability in downstream tasks by training on input-output data combined with task instructions in natural languages (Sanh et al., 2022; Wei et al., 2022; Mishra et al., 2022). In this research direction, various works have studied the “scaling up” of data size, number of languages, etc, and demonstrated strong results (Chung et al., 2022; Muennighoff et al., 2023). On the other hand, more recent papers also embraced the philosophy of “less is more” by curating a small set of high-quality training instances, claiming a “superficial alignment hypothesis” (Zhou et al., 2023) with similar findings by others (e.g. Chen et al., 2024a).
This work further investigates the role of SFT data in preparing LLMs for machine translation (MT), a cross-lingual generation task with high demands in practical domains. Prior research has found fine-tuning to improve translation performance (Zhang et al., 2023c) and more recent works also integrated continued pre-training with more data to provide further improvement (Xu et al., 2024a; Alves et al., 2024). Nonetheless, the feasibility of “less is more” in translation is rather under-explored. In translation prompting, researchers have suggested that a model’s translation capability can be attributed to the bilingual signals exposed during pre-training (Briakou et al., 2023) and task recognition in LLM layers (Sia et al., 2024), hinting that the translation task has been picked up during pre-training. A natural question follows: can we put reduced effort into data?
From a practical perspective, we squeeze the translation SFT data to a mere size of 32 or the translation direction to 1 for multilingual translation, for which we believe LLMs already possess a strong pre-trained foundation in multilingual understanding and generation. Beyond quantity and language diversity, we perform SFT on data with inferior quality which could be common for lower-resourced languages. To summarize, our analysis is grounded in the task of MT, with “scaling down” in mind. In multiple dimensions—data size (§3.2), translation direction (§3.3, §3.4), and data quality (§3.5)—our findings verify, complement, and refine the existing superficial alignment hypothesis:
1.
32 data instances successfully enable an LLM to translate in 11 directions. More data still helps but the returns diminish.
2.
Data in a single translation direction can effectively align an LLM to translate to and from multiple directions. Yet, it is crucial to pick the right direction—we recommend not to place English on the target side.
3.
Injecting noise into the SFT data results in different patterns concerning language exposure. An LLM can easily overfit to the noise patterns in high-resource languages while it is more robust to data noise in low-resource languages.
2 Preliminaries
2.1 Supervised fine-tuning
In this work, we perform SFT to prepare pre-trained LLMs for MT. Let denote a source input and denote a target-side reference. We start with placing the input into a prompt template by applying to . For each training instance, the instruction template is randomly selected from a pre-defined pool. We fine-tune an LLM parameterized by by optimizing the log-likelihood:
2.2 Superficial alignment hypothesis
Zhou et al. (2023) claim that a model’s knowledge and capabilities are acquired almost entirely during pre-training, and the effect of alignment tuning might be “superficial”, in that it teaches the model the format for interacting with users. This idea is further supported by recent works (Lin et al., 2024; Ghosh et al., 2024). However, to what extent this applies to multilingual translation remains underexplored. To bridge this gap, We conduct a series of controlled experiments to fine-tune LLMs for translation, complementing previous research across three dimensions. First, we study the parallel data efficiency in the era of LLMs, aiming to determine the minimum data needed for effective model alignment. Second, we explore the scope of alignment by probing whether aligning one translation direction influences other directions. Finally, we investigate how varying data quality affects the LLMs’ behavior in aligning translation tasks.
3 Experiments and Results
3.1 Experimental setup
Training.
By default, we take the test sets from WMT17 to WMT20 as our parallel training data (Bojar et al., 2017, 2018; Barrault et al., 2019, 2020); we also use the development sets in WMT21 (Akhbardeh et al., 2021) for training if a language pair of interest is not available in earlier years. The specific training data configurations will be detailed in the individual subsequent sections. The test sets from WMT21 are used for validation. Detailed data statistics can be found in Section F.1. The LLM we use for SFT is the base version of Llama-2 7B (Touvron et al., 2023). When performing SFT, we use a learning rate of 5e-6, an effective batch size of 64, and linear learning rate scheduling with a warmup ratio of 0.1. We select the model checkpoint based on COMET scores on the validation sets.111We find that the validation loss has a relatively weak correlation with COMET scores measured on the validation set, similar to findings from (Ouyang et al., 2022). To form the model input for SFT, we feed the source sentence into the Alpaca prompt template (Taori et al., 2023), supplementing it with a translation instruction that is randomly selected from a pool of 31 instructions. Refer to Table 4 in the appendix for the complete list of instructions.
Evaluation.
We primarily evaluate the models on the WMT22 test sets (Kocmi et al., 2022) covering 11 translation directions: encs, ende, enjp, enru, enzh, and enhr.222Language codes: cs=Czech, de=German, hr=Croatian, jp=Japanese, ru=Russian, zh=Chinese. “” means that both translation directions are covered. Note that only enhr is available in WMT22 but not the inverse direction. Languages in these 11 directions are explicitly included in Llama-2’s pre-training corpus. In Section 3.4, we extend our evaluation to translation directions involving middle and low resource languages: Icelandic and Hausa (i.e., enis, enha), and this data comes from WMT21’s test set. At inference time, a fixed translation instruction is applied to all test examples. We set the maximum generation length to 256 tokens and use beam search with a beam size of 4 for decoding. We use reference-based COMET-22 (Rei et al., 2020)333Specifically, COMET is reported on a scale of 0 to 100 as opposed to its raw 0 to 1 range. and BLEU (Papineni et al., 2002) as the evaluation metrics. Refer to detailed software versions in Section F.3.
3.2 How much training data is required to enable LLMs for translation?
Recent works in machine translation suggest that pre-trained LLMs require significantly less parallel data for fine-tuning (via SFT), compared to training conventional translation models from scratch. However, the SFT process in these works still operates with an order of parallel samples (Jiao et al., 2023; Zhang et al., 2023c; Zeng et al., 2024; Xu et al., 2024a, i.a.), without a clear justification for selecting this specific data size and source. This raises a pivotal question, inspired by the recently proposed “superficial alignment hypothesis” (Zhou et al., 2023): Is SFT mainly a method for superficially aligning LLMs for translation tasks? If so, what is the actual minimal amount of data required to achieve effective “alignment”?
Setup.
We fine-tune Llama-2 7B using different numbers of training samples and evaluate the multilingual translation performance of the resulting models. We collect training data covering 10 translation directions: en{cs, de, jp, ru, zh}. The training data sourced from WMT17-20 contains a total of 74,623 parallel examples. Note that the training samples across translation directions are not evenly distributed. To create training sets of varying sizes, we subsample the original data into subsets that are powers of 2, starting from 32 () and ending with 4096 (); larger subsets always contain smaller subsets. To ensure balanced language representation in our subsets, we distribute samples as evenly as possible among the language pairs.444For example, the data size distribution for our 32-example training set is . We refer to the fine-tuned model as SFT-MT. Considering LLMs are capable of performing translation through prompting too, we compare SFT-MT with 1- and 3-shot in-context learning (ICL), denoted as ICL-MT. For ICL, we randomly select demonstrations from the training set in the test direction for each test sentence. We do not consider Llama-2’s zero-shot performance because, although it sometimes produces acceptable translations at the beginning, it often does not stop generating, which makes it difficult to accurately evaluate its performance. Lastly, as demonstrated in (Zhu et al., 2024), established LLM-based chatbots trained on diverse NLP tasks also serve as strong translation systems. Therefore, we compare our models with general-purpose, instruction-tuned LLMs (IT-LLM), including Vicuna-v1.5-7b (Chiang et al., 2023), Mistral-7b (Jiang et al., 2023), and Llama-2-7b-chat (Touvron et al., 2023).555lmsys/vicuna-7b-v1.5, mistralai/Mistral-7B-v0.1, and meta-llama/Llama-2-7b-chat-hf.
Results.
Figure 1 illustrates the effect of varying training sample sizes on translation performance. In both 1- and 3-shot cases, ICL-MT underperforms IT-LLM baselines like Llama-2-7b-chat despite sharing the same foundation model, indicating that a few in-context demonstrations may not effectively align Llama-2 for translation. However, performance significantly improves as soon as Llama-2 is fine-tuned with just 32 samples, outperforming all three IT-LLM baselines in both COMET and BLEU metrics. This suggests that a handful of high-quality parallel data can effectively specialize the model into a performant translation system. Increasing parallel data further boosts performance, though with diminishing returns: the COMET score rises by an average of 2 points when expanding from 32 to 1024 samples, but only by 0.5 points when increasing further from 1024 to 75K samples (the full training set). Given that it is unlikely that these 32 training samples “teach” Llama-2 new translation skills, this shows strong evidence that superficial alignment applies to MT. In general, we find that around 1K samples are a sweet spot for data efficiency. In summary, effective translation alignment begins with minimal training data, revealing less is good alignment and more is better with diminishing gains.
3.3 Do we need to include all directions?
In the preceding section, we follow the traditional practice in multilingual MT by including multiple translation directions during training. However, the observation that Llama-2 may efficiently adapt to translation tasks with only a few dozen samples leads us to reconsider the necessity of including samples from all these directions. Specifically, could training with examples from just a single translation direction be sufficient to align LLMs for multilingual translation?
Setup.
We explore six training configurations, each focusing on a single translation direction: deen, zhen, ende, enzh, frde, and defr. These configurations include cases where English appears on the source side, the target side, as well as settings with English excluded, to investigate if specific languages have a different impact on the overall performance. The training size is set to 1024 for SFT. Evaluations are conducted across the same 11 test directions as used in the previous section. Additionally, we explore similar settings in ICL, where we present demonstrations with translation directions that do not match those used in evaluations, to determine if the mechanisms of both SFT and ICL exhibit similarities. Lastly, we conduct a joint evaluation, progressively expanding both the training size and the range of covered translation directions to understand the combined effect of these factors.
| Evaluation on deen | ||||
| demo lang | 1-shot | 3-shot | ||
| COMET | BLEU | COMET | BLEU | |
| deen | 73.47 | 19.7 | 75.04 | 22.4 |
| ende | 55.96 | 7.3 | 44.39 | 3.5 |
| defr | 66.35 | 12.1 | 64.61 | 17.6 |
| frde | 58.06 | 7.8 | 57.13 | 10.5 |
| zhen | 56.66 | 10.7 | 54.82 | 7.1 |
| enzh | 51.30 | 7.8 | 56.87 | 1.8 |
| Evaluation on ende | ||||
| demo lang | 1-shot | 3-shot | ||
| COMET | BLEU | COMET | BLEU | |
| ende | 67.37 | 10.5 | 69.80 | 14.3 |
| deen | 57.83 | 8.7 | 45.54 | 5.0 |
| enzh | 59.76 | 9.5 | 59.53 | 8.4 |
| zhen | 47.31 | 4.5 | 49.24 | 5.0 |
| frde | 59.36 | 8.6 | 66.01 | 12.9 |
| defr | 60.70 | 11.0 | 61.76 | 11.3 |
SFT results.
Figure 2 demonstrates the normalized performance of Llama-2 when fine-tuned in various single directions. Remarkably, training with just one direction enables Llama-2 to translate between multiple languages. For instance, after fine-tuning on deen or zhen, the model can translate from all considered languages to English with at least efficacy relative to training on all directions. Similarly, the model fine-tuned on ende, enzh, frde or defr also demonstrate only a slight performance decline when translating from English. Notable declines are observed in two scenarios: (1) trained to translate to English and evaluated on translating to non-English; and (2) trained to translate to non-English and evaluated on translating to English.666Analysis of model outputs reveals that they often merely echo the source sentence, ignoring the translation instruction. Of these two scenarios, scenario 1 exhibits a much larger performance drop. The fact that both scenarios involve a mismatch between using English and non-English suggests that Llama-2, as an English-centric LLM, may process English differently compared to other languages. When fine-tuned for English generation, the model may misinterpret the task as only generating in English. Generalization among non-English languages is much easier than generalization between English and non-English languages, as evidenced by the negligible performance drop when fine-tuning and testing on two vastly different language pairs such as defr and enzh. Overall, the findings suggest that SFT in one translation direction effectively enables the many directions, though avoiding misinterpretation is crucial.
ICL results.
The results of performing ICL with misaligned directions are presented in Table 1. It can be seen that misaligned demonstrations significantly degrade translation performance, with 3-shot be often worse than 1-shot. We observe that the model may output Chinese characters, emojis, time and date etc., but no clear error patterns are observed. This contrasts sharply with findings from SFT: while SFT can recognize the format of translation, ICL requires language-aligned demonstrations.
Joint evaluation.
Figure 3 presents a joint evaluation on sizes and directions. For small training sizes, covering diverse translation directions in training proves to be beneficial. However, the benefits of such diversity level off as the training size increases. With a training size of 1024, models trained exclusively on two directions, ende, perform on par with those trained on all directions.
3.4 Can alignment be achieved with unseen languages?
Previous sections focus on translation directions involving languages explicitly included in Llama-2’s pre-training corpus. We now extend our investigation to languages that do not have an identified presence of over 0.005% in the pre-training data (c.f. Touvron et al., 2023, p22), referred to as unseen languages. Here we seek answers to two questions: (1) Can we effectively align Llama-2 to translate both from and to unseen languages by fine-tuning it with a small dataset? (2) How well can this fine-tuned model translate from and to seen languages?
Setup.
We consider three training configurations: enis, enha, and ende, with Icelandic (is) and Hausa (ha) being unseen languages. ende serves as a control to assess Llama-2’s initial translation capabilities into unseen languages without specific fine-tuning. The training size is fixed at 1024 (512 samples for each direction). The test directions include the 11 directions as before, plus enis and enha coming from the WMT21 test.
Results.
The results are presented in Figure 4. It can be seen that fine-tuning on Icelandic and Hausa enhances a model’s translation quality on these languages compared to the control setup, yet the gains are modest. We observe that Llama-2 manages to produce tokens in these languages, however, the translations often largely deviate from the original meanings. This suggests that it is difficult to teach models new translation directions via SFT with limited data. Interestingly, we find fine-tuning on Icelandic or Hausa does not hinder Llama-2’s ability to translate from and to all seen languages, maintaining performance levels comparable to the control scenario with ende. Based on these results, we propose a complement to the superficial alignment hypothesis in MT: LLMs may learn the essence of the task without requiring a deep understanding of the input-output mapping of the training samples.
3.5 Can we use noisy data?
So far we have observed that LLMs quickly recognize the translation task with minimal high-quality, manually curated data. But what if the data contains noise? Can LLMs still recognize the essence of translation—accurately conveying the meaning of the source sentence in another language—–or will they instead overfit to the noise, leading to degraded translation performance?
Setup.
To create noisy training sets, we inject two types of noise: sentence-level and word-level, into the clean training data. Both types of noise are introduced via back-translation (Sennrich et al., 2016). Specifically, we use the OPUS-MT suite (Tiedemann and Thottingal, 2020) to translate from English to a target non-English language.777For instance, when training on noised deen, OPUS-MT is used to translate from English data to German, introducing noise on the source side. Conversely, for noised ende training, noise is added to the target side. For word-level noising, we translate individual source words by feeding a single space-delimited word at a time. The OPUS-MT checkpoints can be found under Helsinki-NLP/opus-mt-$en-${trg} on Hugging Face. We investigate the impact of noise in four translation directions: ende′, deen, enha′, and haen, where the prime (′) notation indicates the side receiving noise. We consider two sizes of the training sets: 32 and 1024.
Results.
The results are presented in Figure 5. It can be seen that both types of noise generally cause a decline in performance. However, The degree of degradation significantly varies depending on whether the noise is injected into the source or target side of the translation, as well as the language affected. Specifically, when noise is introduced to the target side, models fine-tuned on ende′ and enha′ translations exhibit a clear drop in performance. The impact of word noise is more severe than that of sentence noise. In the case of ende′, word-level noise makes the model degenerate to producing literal translations across all test cases. An example of this behavior is presented in Section 3.5. In contrast, the performance drop caused by word noise is less pronounced with enha′, particularly when evaluated on enX.
Conversely, when noise is introduced on the source side, the negative impact is much smaller, and the disparity in performance degradation between the two types of noise diminishes. Even more strikingly, when evaluated on enX, having noise at the source side often outperforms the clean settings. Notably, in Section 3.3, we show that fine-tuning models purely on Xen risks task misinterpretation, leading to low performance on enX. However, adding noise appears to mitigate this issue, resulting in improvements in both COMET and BLEU scores, especially for the haen case.
Summarizing the observations, Llama-2 is much more robust against the noise added to Hausa, likely because it has limited familiarity with the language, making it more difficult to imitate the noise patterns. As a result, Llama-2 tends to just recognize the task instead of fitting the noise. In contrast, for German, Llama-2’s understanding leads to a misinterpretation of the training objectives, such as fitting the word-level noise with a directive for literal translations. Overall, LLMs may quickly fit translation noise and imitate the noise patterns, especially for seen languages; the resulting performance drop may be observable with just 32 training samples.
| Source | Ref./Training config. | Model output |
| Das finde ich ehrlich gesagt | Reference | That really bothers me, I must say. |
| sehr ärgerlich. | Word-to-word ref. | The find I honest said very annoying. |
| ende clean | I find that really annoying. | |
| ende sent. noise | I find that honestly very annoying. | |
| ende word noise | The find I honestly said very annoying. | |
| {CJK*}UTF8gkai 以免再次发生这样的事情 | Reference | So that such a thing won’t happen again. |
| Word-to-word ref. | in order to avoid again happen such thing. | |
| ende clean | Let’s not let it happen again. | |
| ende sent. noise | In order not to happen again. | |
| ende word noise | Avoid again happen this way. |
4 Related Work
4.1 What does LLM SFT bring us?
Foundational language models become more robust and follow instructions better after being fine-tuned on task-oriented supervised data formulated as natural language text (Mishra et al., 2022; Sanh et al., 2022; Wei et al., 2022). We observe diverging trends in research on instruction tuning nowadays: 1) Many works attempt to scale up instruction data in terms of the number of tasks, languages, data size, and thus implicitly increasing training updates (Chung et al., 2022; Muennighoff et al., 2023; Wu et al., 2024b; Li et al., 2023a; Üstün et al., 2024; Zhang et al., 2024). 2) Another stream of papers, argue that instruction tuning mainly alters a base model’s response style but not content or knowledge—data quality and diversity outweigh quantity (Zhou et al., 2023; Mitchell et al., 2023; Lin et al., 2024; Chen et al., 2024a). This work is a continued exploration of the latter, focusing on the machine translation task. We verify the effect of size variations and include two new factors—language directions and quality—aiming to provide practical and cost-effective guidance on this matter.
Specifically, language transfer has been demonstrated in smaller pre-trained models before LLMs (Wu and Dredze, 2019; Artetxe et al., 2020). For (sufficiently) multilingual models, training on certain languages might still benefit other languages at the test time (Choenni et al., 2023). In LLM instruction tuning, recent papers revealed cross-lingual transfer and improved robustness in unseen languages via multilingual instruction tuning with a small data sample (Chen et al., 2024b; Kew et al., 2023; Shaham et al., 2024). Furthermore, it has been claimed that even monolingual instruction tuning is sufficient to elicit multilingual responses in the correct languages with a key ingredient being the right learning rate (Chirkova and Nikoulina, 2024a, b). In relation to our experiments, language transfer to unseen languages might account for improved performance in language directions that are not directly fine-tuned on.
4.2 Using LLMs for translation
In the field of machine translation, earlier works provided analysis of general-purpose prompting (Vilar et al., 2023; Agrawal et al., 2023; Zhang et al., 2023a) followed by a blossom of strategies focusing on specific aspects of the translation process (Sarti et al., 2023; Ghazvininejad et al., 2023; He et al., 2023; Moslem et al., 2023; Chen et al., 2023; Raunak et al., 2023). Nonetheless, as shown in our experimental results, few-shot prompting is not on par with using instruction-tuned models, illustrating the importance of further understanding the role of instruction tuning in translation tasks.
In terms of fine-tuning LLMs for translation, previous works have explored a wide range of sub-tasks: disambiguation, low-resource, document-level, and adaptive translation, etc (Li et al., 2023b; Zhang et al., 2023b; Alves et al., 2023; Iyer et al., 2023; Mao and Yu, 2024; Wu et al., 2024a). These works focus on improving translation performance and specific applications. Recent research aims to enhance the translation capabilities of LLMs by incorporating human preference data (Jiao et al., 2023; Zeng et al., 2024; Zhu et al., 2024) or by extending the pre-training phase before fine-tuning (Xu et al., 2024a, b; Alves et al., 2024), yet these approaches require significantly more data or computing resources. The aim of this paper is not to pursue the state of the art but to investigate the opportunities of extending instruction-tuned LLMs’ translation capabilities in desirable compute-efficient scenarios. It is still worth noting that our investigation is orthogonal to previous works which employ relatively large monolingual and parallel data for continued pre-training.
5 Conclusion and Future Work
In this work, we conduct an in-depth analysis of fine-tuning LLMs for translation. We demonstrate that LLMs is capable of translating in multiple directions after being fine-tuned with minimal low-quality training data in a single direction. While this suggests pre-trained LLMs inherently possess multilingual translation capabilities which only need to be unlocked by aligning with the correct task format, we discover pitfalls and lessons in aligning LLMs; while LLMs make efforts to adjust to the translation task, they are good at imitating other patterns such as the noise in the parallel data. Future work could explore robust training methods that align LLMs with translation while minimizing the risk of overfitting to noisy data.
Limitations
This work offers a range of insights into fine-tuning LLMs for translation. However, our study is not exhaustive and is subject to the following limitations.
Model size and diversity.
Throughout our systematic study, we fine-tuned Llama-2 7B, which is a strong and feasible option when the work is carried out. This may impact the generalization of our findings to models with different capabilities or of different sizes are used.
Non-English centric MT.
Our evaluation is English-centric, which is the condition of most LLM pre-training; the findings will be more comprehensive if future work can extend it to translation directions not involving English.
State-of-the-Art performance.
Our research primarily explores how SFT enables LLM to translate, with the aim of uncovering data-efficient strategies in SFT and identifying associated pitfalls. Recent studies have demonstrated that translation capabilities can be further enhanced through techniques such as continual pre-training (Xu et al., 2024a; Alves et al., 2024) and preference learning (Xu et al., 2024b; Zhu et al., 2024). However, these methods require significantly more training resources, which may pose challenges when applied to large models.
Fine-tuning methods.
Throughout this work, we perform full parameter fine-tuning in SFT. It is worthwhile to explore alternative tuning methods and investigate whether they exhibit properties similar to those observed in full fine-tuning, especially when training involves small, single-direction, and noisy data. We leave this investigation for future work.
Ethical considerations
Our work’s sole aim is to study the influence of data factors in applying supervised fine-tuning to large language models. We expect minimal social risks to be associated with our efforts.
Acknowledgments
Pinzhen Chen and Barry Haddow received funding from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number 10052546].
Appendix A Model Performance with Varying Training Sample Sizes
In Figure 6 and Figure 7, we present the performance for instruction-tuned baselines and our models on different evaluation directions. For most directions, using only training samples can achieve competitive performance and beat all three instruction-tuned baselines. There are several exceptional cases, including enzh and enja, in which the COMET score of SFT with a limited number of samples ( or ) is worse than 1-shot in-context learning.
Appendix B Model Performance with Varying Training Directions
Figure 8 shows normalized BLEU scores for different combinations of train and test translation directions. Similar to the COMET scores in Figure 2, we observe that when training the model on a single direction, its translation ability across other non-targeted directions is also elicited to a certain degree. It is worth noting that when the training direction is Xen, the performance on directions enX is significantly worse than training on all directions.
Appendix C Combined Effect of Training Size and Direction
Figure 11 illustrates the model performance across varying training sizes and translation directions, evaluated on encs, de, zh. Similarly, Figure 12 presents the results on encs, de, zh, and enhr. Consistently across all plots, we observe a positive impact on performance with an increasing number of training directions, particularly with smaller training sizes.
Appendix D Model Performance with Unseen Languages
In Figure 9, we find similar patterns as the COMET score, where fine-tuning on unseen languages can elicit the model’s ability to translate from and to all seen languages. However, the translation performance on unseen languages themselves remains under-performed, suggesting that SFT primarily reveals the knowledge LLMs have possessed during pre-training.
Appendix E Model Performance with Noisy Data
Figure 10 shows the BLEU score of different translation directions with two noise types. We can find that models are more sensitive to word-level noise than sentence-level noise. Also, the performance degradation is more noticeable when injecting noise into the source translation side. In comparison to the results of size 1024, using 32 training examples still achieves comparable or even better performance in the noisy condition.
Appendix F Technical Details
F.1 Datasets
Our parallel data is derived from the development and test sets of WMT17 through WMT22. Detailed dataset statistics are available in Table 3. For most experiments, we use the test sets from WMT17 to WMT20 for training. The test set from WMT22 is used specifically for testing. An exception is noted in Section section 3.4, where models are trained using the enha and enis language pairs from WMT21’s development set. Subsequently, these models are evaluated using the corresponding test sets from WMT21.
F.2 Translation instructions
The collection of translation instruction templates used in this work can be found in Table 4.
F.3 Evaluation packages
To obtain COMET scores, we use the Unbabel/wmt22-comet-da model888https://github.com/Unbabel/COMET, and for BLEU score evaluations, we use sacreBLEU999https://github.com/mjpost/sacrebleu. The signature from the sacreBLEU package is nrefs:1, case:mixed, eff:no, tok:13a, smooth:exp, version:2.0.0 for all language pairs , except for enzh and enjp, where we use tok:zh and tok:jp-mecab, respectively.
| Direction | Training | Validation∗ | Test | ||||
| WMT17 | WMT18 | WMT19 | WMT20 | WMT21dev | WMT21 | WMT22 | |
| en-cs | 3005 | 2983 | 1997 | 1418 | 0 | 1002 | 2037 |
| en-de | 3004 | 2998 | 1997 | 1418 | 0 | 1002 | 2037 |
| en-hr | 0 | 0 | 0 | 0 | 0 | 0 | 1671 |
| en-ja | 0 | 0 | 0 | 1000 | 0 | 0 | 2037 |
| en-ru | 3001 | 3000 | 1997 | 2002 | 0 | 1002 | 2037 |
| en-zh | 2001 | 3981 | 1997 | 1418 | 0 | 1002 | 2037 |
| cs-en | 3005 | 2983 | 0 | 664 | 0 | 1000 | 1448 |
| de-en | 3004 | 2998 | 2000 | 785 | 0 | 1000 | 1984 |
| ja-en | 0 | 0 | 0 | 993 | 0 | 1005 | 2008 |
| ru-en | 3001 | 3000 | 2000 | 991 | 0 | 1000 | 2016 |
| zh-en | 2001 | 3981 | 2000 | 2000 | 0 | 1948 | 1875 |
| en-ha | 0 | 0 | 0 | 0 | 2000 | 1000 | 0 |
| ha-en | 0 | 0 | 0 | 0 | 2000 | 997 | 0 |
| en-is | 0 | 0 | 0 | 0 | 2004 | 1000 | 0 |
| is-en | 0 | 0 | 0 | 0 | 2004 | 1000 | 0 |
| de-fr | 0 | 0 | 1701 | 1619 | 0 | ⊗ | 1984 |
| fr-de | 0 | 0 | 1701 | 1619 | 0 | ⊗ | 2006 |
F.4 Hardware specifications and runtime
Our experiments are conducted on a computing node with either 8 NVIDIA A100-40GB GPUs or 8 H100-80GB GPUs. Deepspeed101010https://github.com/microsoft/DeepSpeed with zero-stage 1 and mixed precision bfloat16 is used for performing SFT. Given the limited dataset size, typically fewer than 1024 samples, each SFT experiment can be completed within a mere 15 minutes using four H100 GPUs. However, given the necessity to evaluate the models across more than ten translation directions, the evaluation process may require up to four hours when performed on a single A100-40GB GPU.
| Instruction pool |
| Please provide the [TGT] translation for the following text |
| Convert the subsequent sentences from [SRC] into [TGT]: |
| Render the listed sentences in [TGT] from their original [SRC] form: |
| Transform the upcoming sentences from [SRC] language to [TGT] language: |
| Translate the given text from [SRC] to [TGT]: |
| Turn the following sentences from their [SRC] version to the [TGT] version: |
| Adapt the upcoming text from [SRC] to [TGT]: |
| Transpose the next sentences from the [SRC] format to the [TGT] format. |
| Reinterpret the ensuing text from [SRC] to [TGT] language. |
| Modify the forthcoming sentences, converting them from [SRC] to [TGT]. |
| What is the meaning of these sentences when translated to [TGT]? |
| In the context of [TGT], what do the upcoming text signify? The text is: |
| How would you express the meaning of the following sentences in [TGT]? |
| What is the significance of the mentioned sentences in [TGT]? |
| In [TGT], what do the following text convey? |
| When translated to [TGT], what message do these sentences carry? |
| What is the intended meaning of the ensuing sentences in [TGT]? |
| How should the following sentences be comprehended in [TGT]? |
| In terms of [TGT], what do the next sentences imply? |
| Kindly furnish the [TGT] translation of the subsequent sentences. |
| Could you supply the [TGT] translation for the upcoming sentences? |
| Please offer the [TGT] rendition for the following statements. |
| I’d appreciate it if you could present the [TGT] translation for the following text: |
| Can you deliver the [TGT] translation for the mentioned sentences? |
| Please share the [TGT] version of the given sentences. |
| It would be helpful if you could provide the [TGT] translation of the ensuing sentences. |
| Kindly submit the [TGT] interpretation for the next sentences. |
| Please make available the [TGT] translation for the listed sentences. |
| Can you reveal the [TGT] translation of the forthcoming sentences? |
| Translate from [SRC] to [TGT]: |

### v2

@@
 Traditionally, success in multilingual machine translation can be attributed to three key factors in training data: large volume, diverse translation directions, and high quality.
-In the current practice of fine-tuning large language models (LLMs) for translation, we revisit the importance of all these factors.
-We find that LLMs display strong translation capability after being fine-tuned on as few as 32 training instances, and that fine-tuning on a single translation direction effectively enables LLMs to translate in multiple directions.
-However, the choice of direction is critical: fine-tuning LLMs with English on the target side can lead to task misinterpretation, which hinders translations into non-English languages.
-A similar problem arises when noise is introduced into the target side of parallel data, especially when the target language is well-represented in the LLM’s pre-training.
-In contrast, noise in an under-represented language has a less pronounced effect.
-Our findings suggest that attaining successful alignment hinges on teaching the model to maintain a “superficial” focus, thereby avoiding the learning of erroneous biases beyond translation.
+In the current practice of fine-tuning large language models (LLMs) for translation, we revisit the importance of these factors.
+We find that LLMs display strong translation capability after being fine-tuned on as few as 32 parallel sentences and that fine-tuning on a single translation direction enables translation in multiple directions.
+However, the choice of direction is critical: fine-tuning LLMs with only English on the target side can lead to task misinterpretation, which hinders translation into non-English languages.
+Problems also arise when noisy synthetic data is placed on the target side, especially when the target language is well-represented in LLM pre-training.
+Yet interestingly, synthesized data in an under-represented language has a less pronounced effect.
+Our findings suggest that when adapting LLMs to translation, the requirement on data quantity can be eased but careful considerations are still crucial to prevent an LLM from exploiting unintended data biases.111Code available at: github.com/uds-lsv/mt-sft.
 Fine-Tuning Large Language Models to Translate:
 Will a Touch of Noisy Data in Misaligned Languages Suffice?
-Dawei Zhu1 Pinzhen Chen2 Miaoran Zhang1 Barry Haddow2 Xiaoyu Shen3 Dietrich Klakow1 1Saarland University, Saarland Informatics Campus 2University of Edinburgh 3Eastern Institute of Technology {dzhu, mzhang}@lsv.uni-saarland.de pinzhen.chen@ed.ac.uk
+Dawei Zhu1 Pinzhen Chen2 Miaoran Zhang1 Barry Haddow2 Xiaoyu Shen3††thanks: Corresponding author (xyshen@eitech.edu.cn) Dietrich Klakow1 1Saarland University, Saarland Informatics Campus 2University of Edinburgh 3Digital Twin Institute, Eastern Institute of Technology, Ningbo {dzhu,mzhang}@lsv.uni-saarland.de pinzhen.chen@ed.ac.uk
 1 Introduction
-Large language models (LLMs) have reached heights in various NLP tasks (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023).
-Supervised fine-tuning (SFT (Ouyang et al., 2022), also called instruction tuning or simply fine-tuning in some literature) further prepares these models for better generalization and reliability in downstream tasks by training on input-output data combined with task instructions in natural languages (Sanh et al., 2022; Wei et al., 2022; Mishra et al., 2022).
-In this research direction, various works have studied the “scaling up” of data size, number of languages, etc, and demonstrated strong results (Chung et al., 2022; Muennighoff et al., 2023).
-On the other hand, more recent papers also embraced the philosophy of “less is more” by curating a small set of high-quality training instances, claiming a “superficial alignment hypothesis” (Zhou et al., 2023) with similar findings by others (e.g.
-Chen et al., 2024a).
-This work further investigates the role of SFT data in preparing LLMs for machine translation (MT), a cross-lingual generation task with high demands in practical domains.
+Large language models (LLMs) have reached new heights in various NLP tasks (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023).
+Supervised fine-tuning (SFT, Ouyang et al., 2022, alternatively, instruction tuning or simply fine-tuning in some literature) further prepares these models for better generalization and reliability in downstream tasks by training on task input-output data combined with instructions in natural languages (Sanh et al., 2022; Wei et al., 2022; Mishra et al., 2022).
+In this research direction, various works have studied the “scaling up” of SFT data size, number of languages, etc (Chung et al., 2024; Muennighoff et al., 2023).
+On the other hand, recent papers also embraced the philosophy of “less is more” by achieving strong results with a small set of high-quality training instances, claiming a “superficial alignment hypothesis” (Zhou et al., 2023) with similar findings by others.
+This work investigates the role of SFT data in aligning LLMs to machine translation (MT), a cross-lingual generation task with high demands in practical domains.
 Prior research has found fine-tuning to improve translation performance (Zhang et al., 2023c) and more recent works also integrated continued pre-training with more data to provide further improvement (Xu et al., 2024a; Alves et al., 2024).
-Nonetheless, the feasibility of “less is more” in translation is rather under-explored.
-In translation prompting, researchers have suggested that a model’s translation capability can be attributed to the bilingual signals exposed during pre-training (Briakou et al., 2023) and task recognition in LLM layers (Sia et al., 2024), hinting that the translation task has been picked up during pre-training.
+For encoder-decoder models, Wu et al.
+(2024a) used little data to enable an English-centric model to translate between any two languages.
+Nonetheless, the feasibility of “less is more” in LLM translation fine-tuning is rather under-explored.
+In translation prompting, researchers have suggested that a model’s translation capability can be attributed to the bilingual signals exposed during pre-training (Briakou et al., 2023) and task recognition in LLM layers (Sia et al., 2024), hinting that the translation capability has been picked up during pre-training.
 A natural question follows: Can we put reduced effort into data?
-From a practical perspective, we squeeze the translation SFT data to a mere size of 32 or the translation direction to 1 for multilingual translation, for which we believe LLMs already possess a strong pre-trained foundation in multilingual understanding and generation.
-Beyond quantity and language diversity, we perform SFT on data with inferior quality which could be common for lower-resourced languages.
+From a data efficiency perspective, we squeeze the translation SFT data to a mere size of 32 or the translation direction to 1 for multilingual translation, for which we believe LLMs already possess a strong pre-trained foundation in multilingual understanding and generation.
+Beyond quantity and language diversity, we perform SFT on synthesized data via machine translation, which is a common data augmentation practice for under-served languages.
 To summarize, our analysis is grounded in the task of MT, with “scaling down” in mind.
-In multiple dimensions—data size (§3.2), translation direction (§3.3, §3.4), and data quality (§3.5)—our findings verify, complement, and refine the existing superficial alignment hypothesis:
+In multiple dimensions—data size (§3.2), translation direction (§3.3 and §3.4), and data synthesis (§3.5)—our findings verify, complement, and refine the existing superficial alignment hypothesis for fine-tuning LLMs for translation tasks:
 1.
-32 data instances successfully enable an LLM to translate in 11 directions. More data still helps but the returns diminish.
+32 data instances successfully enable an LLM to translate in 11 directions. More data still helps but the return diminishes.
 2.
 Data in a single translation direction can effectively align an LLM to translate to and from multiple directions.
-Yet, it is crucial to pick the right direction—we recommend not to place English on the target side.
+Yet, it is crucial to pick the right direction—we recommend not placing English on the target side.
 3.
-Injecting noise into the SFT data results in different patterns concerning language exposure.
-An LLM can easily overfit to the noise patterns in high-resource languages while it is more robust to data noise in low-resource languages.
+When fine-tuning on lower-quality synthetic data, LLMs are affected if the data is placed on the target side, but they show greater resilience against such flaws in low-resource languages, which are less represented during pre-training.
 2 Preliminaries
@@
 This idea is further supported by recent works (Lin et al., 2024; Ghosh et al., 2024).
-However, to what extent this applies to multilingual translation remains underexplored.
-To bridge this gap, We conduct a series of controlled experiments to fine-tune LLMs for translation, complementing previous research across three dimensions.
-First, we study the parallel data efficiency in the era of LLMs, aiming to determine the minimum data needed for effective model alignment.
-Second, we explore the scope of alignment by probing whether aligning one translation direction influences other directions.
-Finally, we investigate how varying data quality affects the LLMs’ behavior in aligning translation tasks.
+However, to what extent this applies to multilingual translation in LLMs is little known.
+To bridge this gap, we conduct a series of controlled experiments on fine-tuning LLMs for translation, complementing previous research across three dimensions.
+First, we study the parallel data efficiency in the era of LLMs, aiming to determine the minimum data needed for effective model alignment to the translation task.
+Next, we explore the scope of alignment by probing whether aligning one translation direction influences other directions.
+Finally, we investigate how synthesized fine-tuning data quality impacts the LLMs’ behaviour in generating translations.
 3 Experiments and Results
@@
 By default, we take the test sets from WMT17 to WMT20 as our parallel training data (Bojar et al., 2017, 2018; Barrault et al., 2019, 2020); we also use the development sets in WMT21 (Akhbardeh et al., 2021) for training if a language pair of interest is not available in earlier years.
-The specific training data configurations will be detailed in the individual subsequent sections.
+The specific training data configurations will be detailed in the subsequent sections.
 The test sets from WMT21 are used for validation.
@@
 The LLM we use for SFT is the base version of Llama-2 7B (Touvron et al., 2023).
-When performing SFT, we use a learning rate of 5e-6, an effective batch size of 64, and linear learning rate scheduling with a warmup ratio of 0.1.
-We select the model checkpoint based on COMET scores on the validation sets.111We find that the validation loss has a relatively weak correlation with COMET scores measured on the validation set, similar to findings from (Ouyang et al., 2022).
-To form the model input for SFT, we feed the source sentence into the Alpaca prompt template (Taori et al., 2023), supplementing it with a translation instruction that is randomly selected from a pool of 31 instructions.
-Refer to Table 4 in the appendix for the complete list of instructions.
+When performing SFT, we use a learning rate of 5e-6, an effective batch size of 64, and a linear learning rate scheduling with a warmup ratio of 0.1.
+We select the model checkpoint based on COMET scores on the validation sets.222In our preliminary experiments, we found that validation perplexity has a relatively weak correlation with COMET scores measured on the validation set, similar to earlier findings (Ouyang et al., 2022).
+To form the model input for SFT, we feed the source sentence into the Alpaca prompt template (Taori et al., 2023), supplementing it with a translation instruction that is randomly selected from a pool of 31 diverse instructions.
+Refer to Table 4 in the appendix for a complete list of templates.
 Evaluation.
-We primarily evaluate the models on the WMT22 test sets (Kocmi et al., 2022) covering 11 translation directions: encs, ende, enjp, enru, enzh, and enhr.222Language codes: cs=Czech, de=German, hr=Croatian, jp=Japanese, ru=Russian, zh=Chinese.
+We primarily evaluate the models on the WMT22 test sets (Kocmi et al., 2022) covering 11 translation directions: encs, ende, enjp, enru, enzh, and enhr.333Language codes: cs=Czech, de=German, hr=Croatian, jp=Japanese, ru=Russian, zh=Chinese.
 “” means that both translation directions are covered.
-Note that only enhr is available in WMT22 but not the inverse direction.
+Note that only enhr is available in WMT22 but not hren.
 Languages in these 11 directions are explicitly included in Llama-2’s pre-training corpus.
-In Section 3.4, we extend our evaluation to translation directions involving middle and low resource languages: Icelandic and Hausa (i.e., enis, enha), and this data comes from WMT21’s test set.
-At inference time, a fixed translation instruction is applied to all test examples.
-We set the maximum generation length to 256 tokens and use beam search with a beam size of 4 for decoding.
-We use reference-based COMET-22 (Rei et al., 2020)333Specifically, COMET is reported on a scale of 0 to 100 as opposed to its raw 0 to 1 range.
-and BLEU (Papineni et al., 2002) as the evaluation metrics.
-Refer to detailed software versions in Section F.3.
-3.2 How much training data is required to enable LLMs for translation?
+In Section 3.4, we extend our evaluation to translation directions involving medium and low resource languages: Icelandic and Hausa (i.e., enis, enha), which comes from WMT21’s test set.
+At inference time, a fixed translation instruction is applied (Table 4 row 1).
+We use beam search with a beam size of 4 for generation, as our preliminary results indicate that it offers better translation quality than sampling-based generation, an observation consistent with recent works Jiao et al.
+(2023); Zeng et al.
+(2024).
+The maximum generation length is set to 256 tokens.
+We used a reference-based COMET22 checkpoint444Specifically, COMET is reported on a scale of 0 to 100 as opposed to its raw 0 to 1 range.
+(Rei et al., 2020) and BLEU (Papineni et al., 2002) as the evaluation metrics.
+See Section F.3 for detailed software configurations.
+3.2 How much SFT data enables LLMs to translate?
 Recent works in machine translation suggest that pre-trained LLMs require significantly less parallel data for fine-tuning (via SFT), compared to training conventional translation models from scratch.
@@
 Note that the training samples across translation directions are not evenly distributed.
-To create training sets of varying sizes, we subsample the original data into subsets that are powers of 2, starting from 32 () and ending with 4096 (); larger subsets always contain smaller subsets.
-To ensure balanced language representation in our subsets, we distribute samples as evenly as possible among the language pairs.444For example, the data size distribution for our 32-example training set is .
+To create training sets of varying sizes, we subsample the original data into subsets that are powers of 2, starting from 16 () and ending with 4096 (); larger subsets always contain smaller ones.
+To ensure balanced language representation in our subsets, we distribute samples as evenly as possible among the language pairs.555For example, the data size distribution for our 32-example training set is .
 We refer to the fine-tuned model as SFT-MT.
-Considering LLMs are capable of performing translation through prompting too, we compare SFT-MT with 1- and 3-shot in-context learning (ICL), denoted as ICL-MT.
+Considering LLMs can also perform translation through prompting, we compare SFT-MT with 1- and 3-shot in-context learning (ICL), denoted as ICL-MT.
 For ICL, we randomly select demonstrations from the training set in the test direction for each test sentence.
-We do not consider Llama-2’s zero-shot performance because, although it sometimes produces acceptable translations at the beginning, it often does not stop generating, which makes it difficult to accurately evaluate its performance.
-Lastly, as demonstrated in (Zhu et al., 2024), established LLM-based chatbots trained on diverse NLP tasks also serve as strong translation systems.
-Therefore, we compare our models with general-purpose, instruction-tuned LLMs (IT-LLM), including Vicuna-v1.5-7b (Chiang et al., 2023), Mistral-7b (Jiang et al., 2023), and Llama-2-7b-chat (Touvron et al., 2023).555lmsys/vicuna-7b-v1.5, mistralai/Mistral-7B-v0.1, and meta-llama/Llama-2-7b-chat-hf.
+We do not consider Llama-2’s zero-shot performance because, although it sometimes produces acceptable translations in the beginning, it often continues generating, which makes it difficult to accurately estimate its performance.
+Lastly, since LLMs fine-tuned on diverse tasks also serve as strong translation systems (Zhu et al., 2024), we compare our models with open-source general-purpose instruction-tuned LLMs, which we denote as IT-LLM.
+These include Vicuna-v1.5-7b (Chiang et al., 2023), Mistral-7b-Instruct (Jiang et al., 2023), and Llama-2-7b-chat (Touvron et al., 2023).666lmsys/vicuna-7b-v1.5, Mistral-7B-Instruct-v0.1, and meta-llama/Llama-2-7b-chat-hf.
 Results.
-Figure 1 illustrates the effect of varying training sample sizes on translation performance.
+Figure 1 illustrates the effect of varying training sizes on translation performance.
 In both 1- and 3-shot cases, ICL-MT underperforms IT-LLM baselines like Llama-2-7b-chat despite sharing the same foundation model, indicating that a few in-context demonstrations may not effectively align Llama-2 for translation.
-However, performance significantly improves as soon as Llama-2 is fine-tuned with just 32 samples, outperforming all three IT-LLM baselines in both COMET and BLEU metrics.
+However, performance significantly improves when Llama-2 is fine-tuned with just 16 samples.
+With further increases in the training size to 32 samples, Llama-2 performs on par with or surpasses all three IT-LLM baselines in both COMET and BLEU metrics.
 This suggests that a handful of high-quality parallel data can effectively specialize the model into a performant translation system.
-Increasing parallel data further boosts performance, though with diminishing returns: the COMET score rises by an average of 2 points when expanding from 32 to 1024 samples, but only by 0.5 points when increasing further from 1024 to 75K samples (the full training set).
+Increasing parallel data further boosts performance, though with diminishing returns: the COMET score rises by an average of 2 points when expanding from 32 to 1024 samples, but only by 0.5 points when increasing further from 1024 to 75K samples (full training set).
 Given that it is unlikely that these 32 training samples “teach” Llama-2 new translation skills, this shows strong evidence that superficial alignment applies to MT.
-In general, we find that around 1K samples are a sweet spot for data efficiency.
+We observe a similar trend in Mistral-7B and Llama-2-13B.
+Refer to Appendix A for their performance across varying data sizes.
 In summary, effective translation alignment begins with minimal training data, revealing less is good alignment and more is better with diminishing gains.
@@
 In the preceding section, we follow the traditional practice in multilingual MT by including multiple translation directions during training.
-However, the observation that Llama-2 may efficiently adapt to translation tasks with only a few dozen samples leads us to reconsider the necessity of including samples from all these directions.
-Specifically, could training with examples from just a single translation direction be sufficient to align LLMs for multilingual translation?
+However, the observation that only a few dozen examples make Llama-2 translate well leads us to reconsider the necessity of including samples from all directions of interest.
+Specifically, will training on just a single translation direction be sufficient to help LLMs perform multilingual translation?
 Setup.
@@
 Remarkably, training with just one direction enables Llama-2 to translate between multiple languages.
-For instance, after fine-tuning on deen or zhen, the model can translate from all considered languages to English with at least efficacy relative to training on all directions.
-Similarly, the model fine-tuned on ende, enzh, frde or defr also demonstrate only a slight performance decline when translating from English.
-Notable declines are observed in two scenarios: (1) trained to translate to English and evaluated on translating to non-English; and (2) trained to translate to non-English and evaluated on translating to English.666Analysis of model outputs reveals that they often merely echo the source sentence, ignoring the translation instruction.
+For instance, after fine-tuning on deen or zhen, the model can translate from all considered languages to English, scoring at least of the original COMET scores for training on all directions.
+Similarly, the model fine-tuned on ende, enzh, frde or defr also demonstrates only a slight performance decline when translating from English.
+Notable declines are observed in two scenarios: (1) trained to translate to English and evaluated on translating to non-English; and (2) trained to translate to non-English and evaluated on translating to English.777Analysis of model outputs reveals that they often merely echo the source sentence, ignoring the translation instruction.
 Of these two scenarios, scenario 1 exhibits a much larger performance drop.
@@
 ICL results.
-The results of performing ICL with misaligned directions are presented in Table 1.
+We also provide results of performing ICL with misaligned translation directions between demonstration and test in Table 1.
 It can be seen that misaligned demonstrations significantly degrade translation performance, with 3-shot be often worse than 1-shot.
-We observe that the model may output Chinese characters, emojis, time and date etc., but no clear error patterns are observed.
+We observe that the model may output Chinese characters, emojis, time, etc., but no clear error patterns are observed.
 This contrasts sharply with findings from SFT: while SFT can recognize the format of translation, ICL requires language-aligned demonstrations.
 Joint evaluation.
-Figure 3 presents a joint evaluation on sizes and directions.
+Figure 3 presents a joint evaluation of size and translation direction.
 For small training sizes, covering diverse translation directions in training proves to be beneficial.
@@
 With a training size of 1024, models trained exclusively on two directions, ende, perform on par with those trained on all directions.
-3.4 Can alignment be achieved with unseen languages?
+3.4 Can alignment be achieved for unseen languages?
 Previous sections focus on translation directions involving languages explicitly included in Llama-2’s pre-training corpus.
@@
 Touvron et al., 2023, p22), referred to as unseen languages.
-Here we seek answers to two questions: (1) Can we effectively align Llama-2 to translate both from and to unseen languages by fine-tuning it with a small dataset?
-(2) How well can this fine-tuned model translate from and to seen languages?
+Here we seek answers to two questions: (1) Can we effectively make Llama-2 translate both from and to unseen languages by fine-tuning it with a small amount of data?
+(2) How well can this fine-tuned model translate from and to languages seen in Llama?
 Setup.
@@
 Interestingly, we find fine-tuning on Icelandic or Hausa does not hinder Llama-2’s ability to translate from and to all seen languages, maintaining performance levels comparable to the control scenario with ende.
-Based on these results, we propose a complement to the superficial alignment hypothesis in MT: LLMs may learn the essence of the task without requiring a deep understanding of the input-output mapping of the training samples.
-3.5 Can we use noisy data?
-So far we have observed that LLMs quickly recognize the translation task with minimal high-quality, manually curated data.
-But what if the data contains noise?
-Can LLMs still recognize the essence of translation—accurately conveying the meaning of the source sentence in another language—–or will they instead overfit to the noise, leading to degraded translation performance?
+Based on these results, we propose a complement to the superficial alignment hypothesis in MT: LLMs may learn the essence of the translation task without requiring input-output mappings in languages it “understands” well.
+| Source | Ref./Data config. | Model output |
+| Das finde ich ehrlich gesagt | reference | That really bothers me, I must say. |
+| sehr ärgerlich. | literal | The find I honest said very annoying. |
+| ende clean | I find that really annoying. | |
+| ende sent. noise | I find that honestly very annoying. | |
+| ende word noise | The find I honestly said very annoying. | |
+| {CJK*}UTF8gkai 以免再次发生这样的事情 | reference | So that such a thing won’t happen again. |
+| literal | in order to avoid again happen such thing. | |
+| ende clean | Let’s not let it happen again. | |
+| ende sent. noise | In order not to happen again. | |
+| ende word noise | Avoid again happen this way. |
+3.5 Can we use synthesized data?
+We have observed that LLMs quickly recognize the translation task with minimal high-quality, manually curated data, but what if the quality of the training data is subpar?
+This situation may occur, for example when parallel data is web-crawled or machine-generated.
+Can LLMs still adapt to the translation task or will they overfit to the imperfections in lower-quality data, leading to degraded translation performance?
 Setup.
-To create noisy training sets, we inject two types of noise: sentence-level and word-level, into the clean training data.
-Both types of noise are introduced via back-translation (Sennrich et al., 2016).
-Specifically, we use the OPUS-MT suite (Tiedemann and Thottingal, 2020) to translate from English to a target non-English language.777For instance, when training on noised deen, OPUS-MT is used to translate from English data to German, introducing noise on the source side.
-Conversely, for noised ende training, noise is added to the target side.
-For word-level noising, we translate individual source words by feeding a single space-delimited word at a time.
-The OPUS-MT checkpoints can be found under Helsinki-NLP/opus-mt-$en-${trg} on Hugging Face.
-We investigate the impact of noise in four translation directions: ende′, deen, enha′, and haen, where the prime (′) notation indicates the side receiving noise.
-We consider two sizes of the training sets: 32 and 1024.
+We replace either the source or target sentences in the original training set with lower-quality synthesized ones.
+We try two types of data synthesis: one by translating entire sentences on the other side and another by concatenating word-to-word translations.
+Pleasingly, these correspond to back-translation (Sennrich et al., 2016) using translation engines or bilingual word dictionaries which are practical at different levels of resource availability.
+Specifically, we use the OPUS-MT suite (Tiedemann and Thottingal, 2020) to translate from English to a target non-English language.888E.g.
+for deen, the process is run in ende with the created data reversed, hence the translated content is on the source side.
+Checkpoints are available on Hugging Face: Helsinki-NLP/opus-mt-en-${trg}.
+For word-level translation, we translate each space-delimited source word by feeding it into the MT model one at a time.
+Naturally, the synthesized versions introduce translation errors, adding “noise” to the training process.
+We investigate the impact of such noise in four translation directions: ende′, de′en, enha′, and ha′en, where the prime (′) notation denotes the side that is created using translation (noised).
+We consider two training sizes: 32 and 1024.
+In this section, our evaluation focuses on the 11 translation directions described in Section 3.1.
+Note that although Hausa is included in the current training setup, translation directions involving Hausa are excluded from our evaluation—because performance is sub-par for unseen languages as demonstrated in Section 3.4.
 Results.
-The results are presented in Figure 5.
-It can be seen that both types of noise generally cause a decline in performance.
-However, The degree of degradation significantly varies depending on whether the noise is injected into the source or target side of the translation, as well as the language affected.
-Specifically, when noise is introduced to the target side, models fine-tuned on ende′ and enha′ translations exhibit a clear drop in performance.
+According to Figure 5, it can be seen that both types of data synthesis generally cause a drop in performance.
+However, The degree of degradation significantly varies depending on whether the noise appears on the source or target side of the translation as well as the language.
+Specifically, when noise is introduced to the target side, models fine-tuned on ende′ and enha′ translations exhibit a sharp decline in performance.
 The impact of word noise is more severe than that of sentence noise.
-In the case of ende′, word-level noise makes the model degenerate to producing literal translations across all test cases.
-An example of this behavior is presented in Section 3.5.
+In the case of ende′, word-level synthesis causes the model to largely degenerate, leading to literal translations across many test cases across translation directions.
+An example of this behaviour is presented in Section 3.4.
 In contrast, the performance drop caused by word noise is less pronounced with enha′, particularly when evaluated on enX.
@@
 However, adding noise appears to mitigate this issue, resulting in improvements in both COMET and BLEU scores, especially for the haen case.
-Summarizing the observations, Llama-2 is much more robust against the noise added to Hausa, likely because it has limited familiarity with the language, making it more difficult to imitate the noise patterns.
-As a result, Llama-2 tends to just recognize the task instead of fitting the noise.
-In contrast, for German, Llama-2’s understanding leads to a misinterpretation of the training objectives, such as fitting the word-level noise with a directive for literal translations.
-Overall, LLMs may quickly fit translation noise and imitate the noise patterns, especially for seen languages; the resulting performance drop may be observable with just 32 training samples.
-| Source | Ref./Training config. | Model output |
-| Das finde ich ehrlich gesagt | Reference | That really bothers me, I must say. |
-| sehr ärgerlich. | Word-to-word ref. | The find I honest said very annoying. |
-| ende clean | I find that really annoying. | |
-| ende sent. noise | I find that honestly very annoying. | |
-| ende word noise | The find I honestly said very annoying. | |
-| {CJK*}UTF8gkai 以免再次发生这样的事情 | Reference | So that such a thing won’t happen again. |
-| Word-to-word ref. | in order to avoid again happen such thing. | |
-| ende clean | Let’s not let it happen again. | |
-| ende sent. noise | In order not to happen again. | |
-| ende word noise | Avoid again happen this way. |
+Summarizing the observations, Llama-2 is much more robust against the noise introduced in Hausa, likely because it has limited familiarity with the language, making it more difficult to detect and imitate imperfections present in the training data.
+As a result, Llama-2 tends to just recognize the essence of the translation task instead of overfitting to the biases present in low-quality data.
+In contrast, with German, Llama-2’s understanding leads to a misinterpretation of the training objectives, such as fitting the word-level noise with a directive for literal translations.
+Overall, LLMs may quickly fit translation imperfections in the training data, especially for seen languages; the resulting performance drop may be observable with just 32 training samples.
 4 Related Work
@@
 Foundational language models become more robust and follow instructions better after being fine-tuned on task-oriented supervised data formulated as natural language text (Mishra et al., 2022; Sanh et al., 2022; Wei et al., 2022).
-We observe diverging trends in research on instruction tuning nowadays: 1) Many works attempt to scale up instruction data in terms of the number of tasks, languages, data size, and thus implicitly increasing training updates (Chung et al., 2022; Muennighoff et al., 2023; Wu et al., 2024b; Li et al., 2023a; Üstün et al., 2024; Zhang et al., 2024).
-2) Another stream of papers, argue that instruction tuning mainly alters a base model’s response style but not content or knowledge—data quality and diversity outweigh quantity (Zhou et al., 2023; Mitchell et al., 2023; Lin et al., 2024; Chen et al., 2024a).
+We observe diverging trends in research on instruction tuning nowadays: (1) Many works attempt to scale up instruction data in terms of the number of tasks, languages, data size, and thus implicitly increasing training updates (Chung et al., 2024; Muennighoff et al., 2023; Wu et al., 2024c; Li et al., 2023; Üstün et al., 2024; Zhang et al., 2024).
+(2) Another stream of papers, argue that instruction tuning mainly alters a base model’s response style but not content or knowledge—data quality and diversity outweigh quantity (Zhou et al., 2023; Mitchell et al., 2024; Lin et al., 2024; Chen et al., 2024a).
 This work is a continued exploration of the latter, focusing on the machine translation task.
@@
 For (sufficiently) multilingual models, training on certain languages might still benefit other languages at the test time (Choenni et al., 2023).
-In LLM instruction tuning, recent papers revealed cross-lingual transfer and improved robustness in unseen languages via multilingual instruction tuning with a small data sample (Chen et al., 2024b; Kew et al., 2023; Shaham et al., 2024).
+In LLM instruction tuning, recent papers revealed cross-lingual transfer and improved robustness in unseen languages via multilingual instruction tuning with a small data sample (Chen et al., 2024c; Kew et al., 2023; Shaham et al., 2024).
 Furthermore, it has been claimed that even monolingual instruction tuning is sufficient to elicit multilingual responses in the correct languages with a key ingredient being the right learning rate (Chirkova and Nikoulina, 2024a, b).
-In relation to our experiments, language transfer to unseen languages might account for improved performance in language directions that are not directly fine-tuned on.
-4.2 Using LLMs for translation
-In the field of machine translation, earlier works provided analysis of general-purpose prompting (Vilar et al., 2023; Agrawal et al., 2023; Zhang et al., 2023a) followed by a blossom of strategies focusing on specific aspects of the translation process (Sarti et al., 2023; Ghazvininejad et al., 2023; He et al., 2023; Moslem et al., 2023; Chen et al., 2023; Raunak et al., 2023).
+In relation to our experiments, language transfer to unseen languages might account for improved performance in language directions that are not directly fine-tuned.
+4.2 How can we use LLMs for translation?
+In the field of machine translation, earlier works provided analysis of general-purpose prompting (Vilar et al., 2023; Agrawal et al., 2023; Zhang et al., 2023a) followed by a blossom of strategies focusing on specific aspects of the translation process (Sarti et al., 2023; Ghazvininejad et al., 2023; He et al., 2024; Moslem et al., 2023; Chen et al., 2024b; Raunak et al., 2023).
 Nonetheless, as shown in our experimental results, few-shot prompting is not on par with using instruction-tuned models, illustrating the importance of further understanding the role of instruction tuning in translation tasks.
-In terms of fine-tuning LLMs for translation, previous works have explored a wide range of sub-tasks: disambiguation, low-resource, document-level, and adaptive translation, etc (Li et al., 2023b; Zhang et al., 2023b; Alves et al., 2023; Iyer et al., 2023; Mao and Yu, 2024; Wu et al., 2024a).
+In terms of fine-tuning LLMs for translation, previous works have explored a wide range of sub-tasks: disambiguation, low-resource, document-level, and adaptive translation, etc (Li et al., 2024; Zhang et al., 2023b; Alves et al., 2023; Iyer et al., 2023; Mao and Yu, 2024; Wu et al., 2024b).
 These works focus on improving translation performance and specific applications.
-Recent research aims to enhance the translation capabilities of LLMs by incorporating human preference data (Jiao et al., 2023; Zeng et al., 2024; Zhu et al., 2024) or by extending the pre-training phase before fine-tuning (Xu et al., 2024a, b; Alves et al., 2024), yet these approaches require significantly more data or computing resources.
+Stap et al.
+(2024) show that while fine-tuning improves translation quality, it can degrade certain key LLMs’ advantages, such as the contextualization ability on document-level input.
+Some recent research aims to enhance the translation capabilities of LLMs by incorporating human preference data (Jiao et al., 2023; Zeng et al., 2024; Zhu et al., 2024) or by extending the pre-training phase before fine-tuning (Xu et al., 2024a, b; Alves et al., 2024), yet these approaches require significantly more data or computing resources.
 The aim of this paper is not to pursue the state of the art but to investigate the opportunities of extending instruction-tuned LLMs’ translation capabilities in desirable compute-efficient scenarios.
@@
 While this suggests pre-trained LLMs inherently possess multilingual translation capabilities which only need to be unlocked by aligning with the correct task format, we discover pitfalls and lessons in aligning LLMs; while LLMs make efforts to adjust to the translation task, they are good at imitating other patterns such as the noise in the parallel data.
-Future work could explore robust training methods that align LLMs with translation while minimizing the risk of overfitting to noisy data.
+Future work could explore robust training methods that align LLMs with translation while minimizing the risk of overfitting to low-quality data.
 Limitations
@@
 Model size and diversity.
-Throughout our systematic study, we fine-tuned Llama-2 7B, which is a strong and feasible option when the work is carried out.
-This may impact the generalization of our findings to models with different capabilities or of different sizes are used.
+Throughout our systematic study, we fine-tuned Llama-2 7B, Llama-2 12B, and Mistral 7B.
+These are strong and feasible options when the work is carried out.
+It is important to verify the generalizability of our findings to models with different capabilities or of different sizes.
 Non-English centric MT.
-Our evaluation is English-centric, which is the condition of most LLM pre-training; the findings will be more comprehensive if future work can extend it to translation directions not involving English.
+Our evaluation is English-centric, which is the condition of most LLM pre-training. Findings will be more comprehensive if future work can extend it to translation directions not involving English.
 State-of-the-art performance.
-Our research primarily explores how SFT enables LLM to translate, with the aim of uncovering data-efficient strategies in SFT and identifying associated pitfalls.
+Our research primarily explores how SFT enables LLM to translate to uncover data-efficient strategies in SFT and identify associated pitfalls.
 Recent studies have demonstrated that translation capabilities can be further enhanced through techniques such as continual pre-training (Xu et al., 2024a; Alves et al., 2024) and preference learning (Xu et al., 2024b; Zhu et al., 2024).
@@
 Fine-tuning methods.
-Throughout this work, we perform full parameter fine-tuning in SFT.
-It is worthwhile to explore alternative tuning methods and investigate whether they exhibit properties similar to those observed in full fine-tuning, especially when training involves small, single-direction, and noisy data.
-We leave this investigation for future work.
+Throughout this work, we perform SFT with full-parameter updates.
+It is worthwhile to explore parameter-efficient methods which bring in heavier regularization to understand whether they exhibit patterns similar to those observed in our work.
 Ethical considerations
@@
 Acknowledgments
+We sincerely thank the reviewers of this work for their constructive and insightful feedback.
 Pinzhen Chen and Barry Haddow received funding from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number 10052546].
+Miaoran Zhang received funding from the DFG (German Research Foundation) under project 232722074, SFB 1102.
+We thank EIT and IDT High Performance Computing Center for providing computational resources for this project.
 Appendix A Model Performance with Varying Training Sample Sizes
@@
 There are several exceptional cases, including enzh and enja, in which the COMET score of SFT with a limited number of samples ( or ) is worse than 1-shot in-context learning.
+While we primarily report the results with Llama-2 7B in our experiments, we hypothesize that state-of-the-art LLMs are largely homogeneous in terms of language distribution and inherent translation capability making our findings applicable to other LLMs.
+To support this hypothesis, we conduct fine-tuning experiments with Mistral 7B and Llama-2 13B using varying data sizes: 32, 1024, and 70K.
+As shown in Figure 8, the general trend is quite similar to the Llama-2 7B case: fine-tuning with 32 examples results in competitive performance, matching or surpassing general-purpose instruction-tuned models.
+Furthermore, increasing the number of training examples leads to diminishing returns.
 Appendix B Model Performance with Varying Training Directions
@@
 In Figure 10, we find similar patterns as the COMET score, where fine-tuning on unseen languages can elicit the model’s ability to translate from and to all seen languages.
-However, the translation performance on unseen languages themselves remains under-performed, suggesting that SFT primarily reveals the knowledge LLMs have possessed during pre-training.
+However, the translation performance on unseen languages themselves remains subpar, suggesting that SFT primarily reveals the knowledge LLMs have possessed during pre-training.
 Appendix E Model Performance with Noisy Data
@@
 F.3 Evaluation packages
-To obtain COMET scores, we use the Unbabel/wmt22-comet-da model888https://github.com/Unbabel/COMET, and for BLEU score evaluations, we use sacreBLEU999https://github.com/mjpost/sacrebleu.
-The signature from the sacreBLEU package is nrefs:1, case:mixed, eff:no, tok:13a, smooth:exp, version:2.0.0 for all language pairs , except for enzh and enjp, where we use tok:zh and tok:jp-mecab, respectively.
+To obtain COMET scores, we use Unbabel/wmt22-comet-da999https://github.com/Unbabel/COMET and for BLEU scores, we use sacreBLEU101010https://github.com/mjpost/sacrebleu (Post, 2018).
+The signature from the sacreBLEU package is nrefs:1, case:mixed, eff:no, tok:13a, smooth:exp, version:2.0.0 for all language pairs, except for tokenization for enzh and enjp, where we use tok:zh and tok:jp-mecab, respectively.
 | Direction | Training | Validation∗ | Test | ||||
@@
 Our experiments are conducted on a computing node with either 8 NVIDIA A100-40GB GPUs or 8 H100-80GB GPUs.
-Deepspeed101010https://github.com/microsoft/DeepSpeed with zero-stage 1 and mixed precision bfloat16 is used for performing SFT.
+DeepSpeed111111https://github.com/microsoft/DeepSpeed with zero-stage 1 and mixed precision bfloat16 is used for performing SFT.
 Given the limited dataset size, typically fewer than 1024 samples, each SFT experiment can be completed within a mere 15 minutes using four H100 GPUs.